Telescope and AGw4 Commissioning at LDG - 11-12 May 2009 MST

Observers:: J. Hill (mtn) , A. Rakich (Tucson)
Telescope Operator: A. Ceranski
Software Support:
Telescope Support: T. Shih, J. Little

Summary

Another photometric night, although we can see a monsoon storm 100 miles to the south in Mexico. Seeing is a little worse than previous nights. We now complain bitterly when the seeing is worse than 1 arcsec. Tonight varied between 0.6 and 1.4 with a few incidents of 2.6.

Verified on-sky that both the GCS patch and the PSF patch work as expected.

Had a few minor telescope/software problems. One of these was that the new ethernet card Jim put in the to-station broke the IDL license. This provided the interesting opportunity for the TO to drive from lbtmu03 for a while. Another was that OSS got hung in a semi-dead state where it was accepting commands, but not reporting back the updated coordinates of the M2. We lost about half an hour total at the beginning. The telescope was quite reliable after we got going.

We believe that GCS active optics is now working well at DG. It is both self-consistent, and in agreement with IDL. The one niggling discrepancy is that IDL measures 50 nm more Z4 (focus) than GCS, and that is a fairly small niggle.

Took several sets of field aberration data with deliberate coma-free mispointing (misalignment) applied to the secondary mirror. This we give a solid empirical calibration of the aberrations and how to correct them. It is also a good way to exercise the system. We also tested John's IRAF script to collect field aberration data automatically. It is lacking compared to Andrew in the sense that it doesn't tell you any kiwi stories along the way.

We took some off-the-y-axis field data at the end of the night, but the seeing was poor. So this was more for exercise of the system than for data that will provide precise measurements. We are discussing whether the Z12-Z21 aberrations help or hurt in poor seeing.

Preparations

Patched versions of GCS and PSF to fix issues we found last night. We are running GCS v2.1.17 and PSF v4.34.

Primary Mirrors have been CO2 cleaned and telescope rebalanced.

S. Hooper has moved the syslog files back to /var/log/messages (at least for tonight).

Details

Open

19:12 Open shutter doors. T=11.6 degC, D=-14.7 degC , Wind ~8 m/sec. Sky is photometric. One thunderhead visible about 100 miles to south in Mexico.

19:19 pointto M5_1246 Tel="ACQUIRE" PAM- gs=0 - failed because rotator not ready

19:20 pointto M5_1246 Tel="ACQUIRE" PAM- gs=0 - failed because Aaron didn't know about selectAGW.

19:21 pointto M5_1246 Tel="ACQUIRE" PAM- gs=0 - succeeded by sky is a little bright still, but shortening the acquire mode exposure time shows that the star is right there. Adjust IE=-44 CA=+9

19:29 pointto M5_1246 Tel="ACQUIRE" PAM- gs=0 - star moves down 2.5 arcsec. Is this the result of unapplied collimation updates from 19:21 to 19:28?

19:33 pointto BS9143 Tel="ACTIVE" PAM+ PA=0 gs=14 - sextractor did a good job of plucking the star out of twilight. seeing is settling into 0.7 arcsec as twilight fades.

19:40 stopWFS so I can work on a Python script to do agw_collimate, but this is delayed because the IDL license manager is down. Fortunately Jim D. is sitting right here in the control room. Unfortunately, the license program seems to have been married to the ethernet card that Jim removed from the to-station this afternoon.

19:59 pointto BS9143 Tel="ACTIVE" PAM+ PA=0 gs=14

20:14 stop Guiding

20:15 Setup Aaron on lbtmu03 while Jim changes the ethernet card on lbtmu01. This cures the IDL license problem. Time lost 15 minutes.

20:27 OSS is not reporting the correct M2 position although there is some evidence that the mirror is actually moving. John homes the mirror but the readouts on OSS GUI are still frozen. Tried to stop OSS several times but it was in a semi-dead state that needed a kill -9 to stop it. Once properly restarted the M2 readouts are OK Lost 10 minutes.

20:40 Just when we get everything running again, the seeing falls apart. Increase Right Vent Doors to full, and Left to midway.

20:45 seeing is back to 0.6 arcsec.

20:48 Wouldn't accept +300 Z22.

20:49 Send Z22 = -300 during wfsc 0528. Forces on M1 changed, but we don't immediately see Z22.

20:51 Clear active optics on both M1 and M2 and put arbitrator in M1 mode rather than split mode.

20:55 Seeing flares to 1.3 arcsec.

20:58 Stop WFSing to run collimation in IDL

20:58 Seeing flares to 2.6 arcsec.

Testing PSF Z22 focus compensation

21:01 Send +300 Z22. The seeing is bad, but the Z22 changed and Z11/Z4 didn't change. We should test it again in better seeing.

Script development while the seeing is bad

20:03 Stop GCS WFSing to run IDL

21:08 First successful test of agw_collimate.py (a wrapper script to call the IDL command).

21:26 First successful test of agw_wfsc_multiple.py (The makings of a field aberration script in IRAF.)

Testing Off-Axis Active Optics

21:29 pointto BS9152 Tel="ACTIVE" PAM+ PA=INDEF=229.72 gs=1 - seeing is 1.1-1.4 arcsec

21:36 Seeing improves to 0.8 arcsec.

21:40 pointto BS9152 Tel="ACTIVE" PAM+ PA=259.72 gs=1 - (moving off the y-axis) Turn off sending Zernikes to PSF starting with wfsc 0063

21:43 Turn on sending Zernikes to PSF starting with wfsc 0067.

21:49 pointto BS9152 Tel="ACTIVE" PAM+ PA=INDEF=229.72 gs=1 and Turn off sending Zernikes to PSF starting with wfsc 0074.

21:53 Turn on sending to PSF with wfsc 0079. Seeing is back to 0.7, but then 1.3.

21:56 Send Z22 = +300 nm. Turn off sending to PSF. But Bad new influence forces.

21:57 Refocus with sending on. The focus difference between IDL and GCS seems to be an offset of ~50 nm rather than a scale factor. IDL is ~50-70 nm more positive.

22:00 Send -300 Z22 while turning off the sending to PSF. This leaves us with a 2000 nm focus error.

22:04 Clear C00

22:07 Send -300 Z22 and turning off sending during wfsc 0096.

22:09 Send +300 Z22 during wfsc 0100

22:12 Turn on sending to PSF for wfsc 0104.

22:14 seeing improves to 0.6.

Field aberration data

Field Aberration Measurements
Using IDL

deliberately introduce +40" of M2 tilt about coma neutral point as a pointing offset, then take field data. M2 total collimation X-> -9.053, Y-> -1.317, Rx->148.650, Ry-> -210.999 (X and Ry affected by tilt offset).

IE, CA original -47, 4; after pointing adjustment in M2 arbitrator; -48, -34

(Note that a +40" tilt in the PSF arbitratio for M2 coma free pointing required a -38" correction in CA, so the sign is right there).

BS9152
GS PA WFSC
0 0 16-17
17 124.77 18-20
0 124.77 21-22
18 65.44 23-25
0 65.44 26
7 10.54 27-29
0 10.54 30-31
4 338.61 32-34
0 338.61 35
1 229.72 36-38
0 229.72 39
9 168.12 40-42 probably 169.12
0 168.12 43-44
17 124.77 45-47
0 124.77 48
18 65.44 49-51
0 65.44 52
7 10.54 53-55
0 10.54 56
4 338.61 57-59
0 338.61 60-61
1 229.72 62-64
0 229.72 65
9 168.12 66-68 probably 169.12
0 168.12 69-70
17 124.77 71-73
0 124.77 74-75
18 65.44 76-78
0 65.44 79-80
7 10.54 81-83
0 10.54 84-85
4 338.61 86-88
0 338.61 89-90
1 229.72 91-93
0 229.72 94
9 168.12 95-97 probably 169.12
0 168.12 98

Now undoing the M2 coma free 40" tilt and putting in 40" tip. Centering IE CA first without any tip or tilt, IE-> -45 CA-> 11 ; After sending "40" tip on M2, IE->-85 , CA-> 12 .

Semi-Automated Field Aberration Measurements

00:38 Tests of John's script to take aberration data - starting with wfsc 0099

00:44 cl < home$field_BS9152.cl - need to do prepdir first for pointto to run locally.

00:46 cl < home$field_BS9152.cl - must be in proper directory on to-station

00:50 cl < home$field_BS9152.cl - command is multiple not measure

01:00 cl < home$field_BS9152.cl - starting with wfsc 0103 Seeing is 1.2 arcsec

Script looks like this. It doesn't generate log files yet, so I still filled in the WFSC numbers in the wiki by hand.
# IRAF script to automatically collect Field Aberration data
# Created by J. Hill 11-MAY-2009
# Must run as LBTO or with aliases to the python wrappers of IDL commands
# Must run on to-station or other machine with TCS and IRAF
# Must run in image directory such as /Repository/AGW_Data/091205w/
# Must run LBTtools.IRTC.prepdir in that image directory
#
#
pointto ("BS9152", Tel="GUIDE", PAMode=yes, PA=0,  gs=0)
!agw_collimate -e 30
!agw_collimate -e 30
!agw_collimate -e 30
!agw_collimate -e 30

pointto ("BS9152", Tel="GUIDE", PAMode=yes, PA=124.77,  gs=17)
!agw_wfsc_multiple -c 3 -e 30
pointto ("BS9152", Tel="GUIDE", PAMode=yes, PA=124.77,  gs=0)
!agw_collimate -e 30
!agw_collimate -e 30

pointto ("BS9152", Tel="GUIDE", PAMode=yes, PA=65.44,  gs=18)
!agw_wfsc_multiple -c 3 -e 30
pointto ("BS9152", Tel="GUIDE", PAMode=yes, PA=65.44,  gs=0)
!agw_collimate -e 30
!agw_collimate -e 30

BS9152
GS PA WFSC
0 0 0103-0106
17 124.77 0107-0109
0 124.77 0110-0111
18 65.44 0112-0114
0 65.44 0115-0116
7 10.54 0117-0119
0 10.54 0120-0121
4 338.61 0122-0124
0 338.61 0125-0126
1 229.72 0127-0129
0 229.72 0130-0131
9 169.12 0132-0134
0 169.12 0135-0136

01:45 cl < home$field_BS9152.cl - starting with wfsc 0137, seeing is 0.8-1 arcsec

0 0.0 0137-0140
17 124.77 0141-0143
0 124.77 0144-0145
18 65.44 0146-0148
0 65.44 0149-0150
7 10.54 0151-0153
0 10.54 0154-0155
4 338.61 0156-0158
0 338.61 0159-0160
1 229.72 0161-0163
0 229.72 0164-0165
9 169.12 0166-0168
0 169.12 0169-0170

02:25 cl < home$field_BS9152.cl - starting with wfsc 0171,

0 0.0 0171-0174
17 124.77 0175-0177
0 124.77 0178-0179
18 65.44 0180-0182
0 65.44 0183-0184
7 10.54 0185-0187
0 10.54 0188-0189
4 338.61 0190-0192
0 338.61 0193-0194
1 229.72 0195-0197
0 229.72 0198-0199
9 169.12 0200-0202
0 169.12 0203-0204

03:04 complete

0 0.0  
17 124.77  
0 124.77  
18 65.44  
0 65.44  
7 10.54  
0 10.54  
4 338.61  
0 338.61  
1 229.72  
0 229.72  
9 169.12  
0 169.12  

More data on off-axis collimation

03:09 ACQUIRE StoneMagw gs=0 Adjust IE=-79 CA=+2

03:15 ACTIVE gs=0 PA=0 - star is offset by ~8 arcsec. Why?

03:17 ACTIVE gs=38 PA=INDEF=343.44

03:21 ACTIVE gs=83 PA=INDEF=96.87

03:24 ACQUIRE StoneMagw gs=0 Adjust IE=-90 CA=0

03:29 cl < home$field_StoneM.cl (using the text of the script as a log entry) This is a different flavor of the script to measure off-axis aberrations being applied by GCS (as opposed to on-the-y-axis field aberrations being measured above). This is a more systematic version of the tests we did by hand at the beginning of the night. Probably it isn't very useful as the seeing has degraded. Maybe we can use it as an estimate of the noise in the Z12-Z21 aberrations induced by seeing.

pointto ("StoneMagw", Tel="ACTIVE", PAMode=yes, PA=343.44, gs=38) - wfscimage 0123-0125
sleep(120)
pointto ("StoneMagw", Tel="GUIDE", PAMode=yes, PA=343.44, gs=0)
agw_wfsc_multiple -c 3 -e 30 - wfsc 0205-0207

pointto ("StoneMagw", Tel="ACTIVE", PAMode=yes, PA=313.44, gs=38) - wfscimage 0126-0128
sleep(120)
pointto ("StoneMagw", Tel="GUIDE", PAMode=yes, PA=313.44, gs=0)
agw_wfsc_multiple -c 3 -e 30 - wfsc 0208-0210

BUT WE FORGOT TO REMOVE ANDREW'S PURPOSEFUL MISALIGNMENT

03:41 Remove it from M2 pointing offset in PSF. Adjust IE=-50 and the guider recovers the star.

03:43 cl < home$field_StoneM.cl - seeing is fluctuating 1 to 1.8 arcsec for this data

pointto ("StoneMagw", Tel="ACTIVE", PAMode=yes, PA=343.44, gs=38) - wfscimage 0129-0131
sleep(120)
pointto ("StoneMagw", Tel="GUIDE", PAMode=yes, PA=343.44, gs=0)
agw_wfsc_multiple -c 3 -e 30 - wfsc 0211-213

pointto ("StoneMagw", Tel="ACTIVE", PAMode=yes, PA=313.44, gs=38) - wfscimage 0132-0134
sleep(120)
pointto ("StoneMagw", Tel="GUIDE", PAMode=yes, PA=313.44, gs=0)
agw_wfsc_multiple -c 3 -e 30 - wfsc 0214-0216

pointto ("StoneMagw", Tel="ACTIVE", PAMode=yes, PA=13.44, gs=38) - wfscimage 0135-0137 probe x: 95.356 y: 447.359
sleep(120)
pointto ("StoneMagw", Tel="GUIDE", PAMode=yes, PA=13.44, gs=0)
agw_wfsc_multiple -c 3 -e 30 - wfsc 0217-0219

pointto ("StoneMagw", Tel="ACTIVE", PAMode=yes, PA=23.44, gs=38) - wfscimage 0138-0139 probe x: 122.789 y: 466.370
sleep(120)
pointto ("StoneMagw", Tel="GUIDE", PAMode=yes, PA=23.44, gs=0)
agw_wfsc_multiple -c 3 -e 30 -wfsc 0220-0222

pointto ("StoneMagw", Tel="ACTIVE", PAMode=yes, PA=303.44, gs=38) - wfscimage 0140-0142
sleep(120)
pointto ("StoneMagw", Tel="GUIDE", PAMode=yes, PA=303.44, gs=0)
agw_wfsc_multiple -c 3 -e 30 - wfsc 0223-0225

# change stars

pointto ("StoneMagw", Tel="ACTIVE", PAMode=yes, PA=96.87, gs=83) - wfscimage 0143-0145 probe x: 0.0 y: 425.056
sleep(120)
pointto ("StoneMagw", Tel="GUIDE", PAMode=yes, PA=96.87, gs=0)
agw_wfsc_multiple -c 3 -e 30 - wfsc 0226-0228

pointto ("StoneMagw", Tel="ACTIVE", PAMode=yes, PA=66.87, gs=83) - wfscimage 0146-0147
sleep(120)
pointto ("StoneMagw", Tel="GUIDE", PAMode=yes, PA=66.87, gs=0)
agw_wfsc_multiple -c 3 -e 30 - wfsc 0229-0231

pointto ("StoneMagw", Tel="ACTIVE", PAMode=yes, PA=126.87, gs=83) - wfscimage 0148-0149 x: 94.137 y: 450.258
sleep(120)
pointto ("StoneMagw", Tel="GUIDE", PAMode=yes, PA=126.87, gs=0)
agw_wfsc_multiple -c 3 -e 30 - wfsc 0232-0234 - Andrew is making high order astigmatism adjustments during these.

04:35 ACTIVE gs=82 PA=INDEF=96.87

Andrew is adjusting Z12-Z21 aberrations while GCS collimates. Finish with 109 nm rms and 1 arcsec seeing.

Close

04:46 Close - T=7.5 degC (more than double the nightly temperature change compared to last night), D=-16.7 Wind ~6 m/sec. Sky clear but not quite photometric.

-- JohnHill - 12 May 2009
Topic revision: r4 - 12 May 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback